OpenStudy (anonymous):

Suppose that an account earns 5.3% interest and after 3 years the balance is 6954$.Find the principle that was invested.Use the formula B=P+Prt,which is used to calculate the balance in an account if simple interest is added to a principle P at an interest rate of r after t years.

OpenStudy (anonymous):

@help123please.

OpenStudy (anonymous):

@hba

OpenStudy (anonymous):

I seriously need help

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

@mathstudent55

OpenStudy (shamil98):

Would you not just plug in the given values? try doing the equation.

OpenStudy (anonymous):

okay 6954=P+P(5.3/100)(3)

OpenStudy (anonymous):

Am i correct so far?

OpenStudy (jdoe0001):

\(\bf Balance = Principal + Principal \times rate \times time\\ B = P + Prt\\ r = \textit{rate in decimals, 5.3\%} = \frac{5.3}{100}=0.053\\ t = \textit{time in years, in this case } \quad 3\\ B = \textit{the balance in the account, well, today's balance is}\quad \$6,954\\ \quad \\ B = P + Prt\implies 6954= p + p \times 0.053 \times 3\) solve for "p"

OpenStudy (shamil98):

6954 = P + (5.3(3)p)

OpenStudy (shamil98):

yeah solve for p, my input was incorrect , 5.3% interest in decimal is actually 0.53

OpenStudy (anonymous):

i got 43735.84....

OpenStudy (anonymous):

is this right

OpenStudy (shamil98):

0.053* my bad ._.

OpenStudy (anonymous):

i still got this answer i know im wrong but i dont know what i did wrong

OpenStudy (anonymous):

hba help

OpenStudy (shamil98):

6954=p+p×0.053×3 6954 = 2p x 0.159 6954 = 0.318p p = 21867.9245283